OpenStudy (toxicsugar22):

calculate the value of [H30+] from the give [-OH] and label the solution as acidic or basic A. 2.00*10^-3 M B. 9.59*10^-9

OpenStudy (joannablackwelder):

[H3O+][OH-]=1x10^-14

OpenStudy (joannablackwelder):

Can you calculate [H3O+] for A?

OpenStudy (toxicsugar22):

walk me through the steps

OpenStudy (toxicsugar22):

how do i do that

OpenStudy (joannablackwelder):

Substitute the value for OH- and solve for H3O+

OpenStudy (joannablackwelder):

You can do it! :)

OpenStudy (toxicsugar22):

nope i tried

OpenStudy (joannablackwelder):

What did you try?

OpenStudy (toxicsugar22):

on my own i dont see a way

OpenStudy (toxicsugar22):

i tried doing -log(2.00*10^-3 M)

OpenStudy (joannablackwelder):

Why are you using that equation?

OpenStudy (joannablackwelder):

That would solve for pOH, but we aren't looking for that now.

OpenStudy (joannablackwelder):

Use the equation I gave you.

OpenStudy (toxicsugar22):

1x10^-14

OpenStudy (toxicsugar22):

i dont know help me and ill folow along

OpenStudy (toxicsugar22):

please

OpenStudy (joannablackwelder):

[H3O+][OH-] = 1x10^-14 Can you substitute the value for [OH-] given?

OpenStudy (toxicsugar22):

im slowly saying no becuae I dont really know

OpenStudy (joannablackwelder):

Ok. I will show you. [H3O+][2x10^-3] = 1x10^-14

OpenStudy (joannablackwelder):

Do you see how I did that?

OpenStudy (toxicsugar22):

yes

OpenStudy (joannablackwelder):

Ok, good. :) Now, can you solve for [H3O+]?

OpenStudy (toxicsugar22):

dont let me do this bymyself please. I am pullingmy hair out just foing this problem im not really savvy in chemistry

OpenStudy (joannablackwelder):

I am helping you. This part is the algebra you did in the last problem. You can do it!

OpenStudy (joannablackwelder):

Think of it like this: (2x10^-3)x=1x10^-14

OpenStudy (joannablackwelder):

Solve for x.

OpenStudy (toxicsugar22):

it it x=1/20000000

OpenStudy (joannablackwelder):

Close. I get 1/200000000000 or 5x10^-12

OpenStudy (toxicsugar22):

so 5x10^-12is the anwer for A

OpenStudy (toxicsugar22):

SOLUTION

OpenStudy (joannablackwelder):

(1x10^-14)/(2x10^-3)

OpenStudy (joannablackwelder):

Right.

OpenStudy (joannablackwelder):

That is the first part for A.

OpenStudy (joannablackwelder):

We still need to find whether the solution is acid or basic though.

OpenStudy (joannablackwelder):

This you can get using the equation you mentioned before but with [H3O+].

OpenStudy (joannablackwelder):

pH=-log[H3O+] and pH of less than 7 is acidic, greater than 7 is basic

OpenStudy (toxicsugar22):

i i meant basic as it is greater than 7

OpenStudy (joannablackwelder):

Yes! I get pH =11.3 which is basic :)

OpenStudy (toxicsugar22):

now we have to do part b

OpenStudy (joannablackwelder):

Ok.It is the same exact procedure. Can you try?

OpenStudy (toxicsugar22):

[H3O+][ 9.59*10^-9] = 1x10^-14

OpenStudy (toxicsugar22):

is it

OpenStudy (joannablackwelder):

Godd!

OpenStudy (joannablackwelder):

*good haha

OpenStudy (toxicsugar22):

(9.59*10^-9)x = 1x10^-14

OpenStudy (joannablackwelder):

Yup

OpenStudy (toxicsugar22):

x=1.04275*10^-6

OpenStudy (joannablackwelder):

Me too!

OpenStudy (toxicsugar22):

and it is acidic

OpenStudy (toxicsugar22):

so it is 1.04*10^-6

OpenStudy (toxicsugar22):

right

OpenStudy (joannablackwelder):

Right! You did it! :)

OpenStudy (toxicsugar22):

and it is acid

OpenStudy (joannablackwelder):

Correct.

OpenStudy (toxicsugar22):

and it 1.04*10^-6 to three sig figs like 1.04 or is it just 1

OpenStudy (joannablackwelder):

1.04 is 3 sig figs

OpenStudy (toxicsugar22):

(2.33*10^-10)x=1x10^-14

OpenStudy (toxicsugar22):

can you help me solve that

OpenStudy (toxicsugar22):

i got 0.0000429185

OpenStudy (toxicsugar22):

right

OpenStudy (joannablackwelder):

Yup!

OpenStudy (joannablackwelder):

Or 4.29x10^-5

OpenStudy (toxicsugar22):

and it is acidic

OpenStudy (joannablackwelder):

Is the 2.33x10^-10 the [H3O+]?

OpenStudy (joannablackwelder):

or [OH-]?

OpenStudy (joannablackwelder):

pH=-log[H3O+]

OpenStudy (toxicsugar22):

NO SRRY IT IS OH-

OpenStudy (joannablackwelder):

Then yup :)

OpenStudy (toxicsugar22):

i meant acidic

OpenStudy (toxicsugar22):

rigt

OpenStudy (joannablackwelder):

Yes, acidic